Roy Travel offers helicopter service from suburban towns to John F. Kennedy International Airport in New York City. Each of its 12 helicopters makes between 1,200 and 1,700 ​round-trips per year. The records indicate that a helicopter that has made 1,200 ​round-trips in the year incurs an average operating cost of ​$900 per​ round-trip, and one that has made 1,700 ​round-trips in the year incurs an average operating cost of ​$700 per​ round-trip.

1.

Using the​ high-low method, estimate the linear relationship y=a+bX​, where y is the total annual operating cost of a helicopter and X is the number of​round-trips it makes to JFK airport during the year.

2.

Give examples of costs that would be included in a and in b.

3.

If Roy Travel expects each helicopter to​ make, on​ average,1,500
​round-trips in the coming​ year, what should its estimated operating budget for the helicopter fleet​ be?
Answer #1

Solution 1:

Variable cost per round trip = (Cost at high level - Cost at low level) / (high level activity - Low level activity)

= (1700*$700 - 1200*$900) / (1700-1200) = $220 per round trip

Fixed cost = 1700 *(700-$220) = $816,000

Linear relation ship = Y = $816,000 + $220 X

Solution 2:

cost that can be included in a = Salaries expense, airport fees, licence fees etc

Cost that can be included in b = Fuel, meals etc

Solution 3:

Estimated operating budget for 1500 round trips = $816,000 + (1500*$220) = $1,146,000